# Dejavu: Towards Experience Feedback Learning for Embodied Intelligence

### v2

@@ Abstract
Embodied agents face a fundamental limitation: once deployed in real-world environments to perform specific tasks, they are unable to acquire additional knowledge to enhance task performance. In this paper, we propose a general post-deployment learning framework Dejavu, which employs an Experience Feedback Network (EFN) and augments the frozen Vision-Language-Action (VLA) policy with retrieved execution memories. EFN identifies contextually prior action experiences and conditions action prediction on this retrieved guidance. We adopt reinforcement learning with semantic similarity rewards to train EFN, ensuring that the predicted actions align with past behaviors under current observations. During deployment, EFN continually enriches its memory with new trajectories, enabling the agent to exhibit “learning from experience”. Experiments across diverse embodied tasks show that EFN improves adaptability, robustness, and success rates over frozen baselines. We provide code and demo in our supplementary material.
@@ 1 Introduction
Embodied intelligence studies agents that learn and act through physical interaction with the environment [Liu2025EmbodiedAI, sun2024comprehensive, liu2025embodied]. Recently, unified Vision–Language–Action (VLA) models have achieved impressive generalization across diverse tasks [Zitkovich2023RT2, survey1, survey2, survey3], but only after massive offline training on fixed data distributions [Brohan2023RT1, Ha2024Octo]. Once deployed, their weights (and thus their knowledge) remain fixed and are not updated without retraining [Liu2024IJRROnTheJob], so most real systems effectively “stop learning” in the wild [Liu2024IJRROnTheJob].
@@
This naturally raises the question: must intelligent systems always rewrite their internal weights to improve? Humans often tackle new problems by recalling and reusing past experiences instead of acquiring entirely new facts [Andrychowicz2017HER, Oh2018SIL]. Episodic memories, which underlie the feeling of “déjà vu”, do not modify core knowledge representations but enable rapid adaptation by analogy [r2a]. We therefore ask whether an AI agent can similarly improve by recalling and reusing its own “experiences”. If a neural network could leverage stored memories of situations and solutions to inform its current decisions, it could become better over time at inference, simply by accumulating and exploiting new experiences without gradient-based retraining [Blundell2016MFEC, Pritzel2017NEC, Kuznetsov2021EMAC]. This idea of “learning from déjà vu” motivates our approach.
@@
A closely related line of work is retrieval-augmented reinforcement learning and retrieval-augmented embodied agents. Retrieval-augmented RL couples an external experience buffer with a parametric policy or value function, retrieving past trajectories similar to the current state to improve credit assignment or value estimation [r2a]. In embodied and robotic control, recent methods retrieve prior executions in various forms (e.g., state–action trajectories or policy snippets) and leverage them during decision-making, either by fine-tuning on the retrieved data [Zhu2024RAEA, Memmel2025STRAP] or by conditioning the policy directly on retrieved trajectories as additional context [Kuroki2024MABR, Kuang2025RAM, Kwon2025RTCache]. These works share a retrieval-plus-decision paradigm [r2a, Zhu2024RAEA, Memmel2025STRAP], but are typically designed to improve a trainable policy whose weights continue to be updated during or after deployment [Kuang2025RAM, Kuroki2024MABR], operate on static offline corpora rather than live memories that grow in deployment [r2a, Zhu2024RAEA], and define retrieval over compact state or task abstractions instead of the rich, open-vocabulary vision–language interface of modern VLAs [Zhu2024RAEA, Kwon2025RTCache]. As a result, existing retrieval-augmented methods do not by themselves provide a simple deployment-time mechanism for continually improving a frozen, unified policy.
@@
To this end, we propose the Experience Feedback Network (EFN), an experience-centric controller for deployment-time improvement of pre-trained VLA policies without modifying their weights. EFN takes the current observation together with an action retrieved from a matched experience and predicts a residual action, which is added to the VLA policy’s output to produce the final control. When a strong prior experience exists, EFN learns to exploit it to refine the action, while reverting toward the base policy when retrieval is uninformative. We train EFN with reinforcement learning using a dense similarity-based reward: the agent receives higher reward when the next observation resembles the next observation in the retrieved experience, providing frequent shaping signals beyond sparse success–failure feedback. EFN is optimized with the soft actor–critic algorithm [Haarnoja2018SAC], whose entropy regularization and off-policy sample efficiency facilitate effective reuse of stored experiences.
@@
During deployment, we maintain a live experience bank that is continuously augmented with newly successful rollouts. At each inference step, the agent retrieves a similar trajectory in a joint vision–language embedding space, provides the matched experience action to EFN alongside the current observation, and obtains a residual that refines the base VLA output. This yields a retrieval-augmented, yet weight-frozen, VLA agent that improves purely by accumulating and reusing its own experiences over time. We integrate EFN with OpenVLA [openvla], UniVLA [univla], and GO-1 [go1], and evaluate on both simulation environments and a real-world platform. Extensive experiments show that EFN not only improves success and efficiency over strong VLA baselines, but also outperforms representative retrieval-augmented RL and embodied baselines under the same experience bank and evaluation protocol. We summarize our main contributions as follows:
@@
  • We propose EFN, a retrieval-conditioned residual module that augments a frozen VLA policy with an episodic experience bank, enabling post-deployment improvement through memory updates rather than backbone finetuning.
@@
  • We design an experience bank that stores synchronized vision–language–action trajectories, retrieves task-relevant transitions in a joint embedding space, and trains the residual with similarity-shaped reinforcement signals toward retrieved successors.
@@
  • We integrate EFN with OpenVLA, UniVLA, and GO-1 and benchmark it on LIBERO and a real-world AgiBot-G1 setup, showing consistent gains over retrieval-only, residual-only, retrieval-augmented RL, and test-time training baselines under the same interaction budget.
@@
Figure 1: EFN trains a residual policy with SAC to nudge the base action so the next frame matches the retrieved memory’s successor.
@@
Figure 2: EFN infers by retrieving efficient candidates, applies the residual correction and grows the experience bank online.
@@ 2.1 Vision-Language-Action (VLA) Models
Large-scale VLA policies couple open-vocabulary perception with end-to-end control and achieve strong cross-task generalization in manipulation [smolvla, vla_survey, wall-oss, gr3, Brohan2023RT1, Zitkovich2023RT2]. Recent open-source generalist policies [Ha2024Octo, openvla] extend this to heterogeneous robots and sensors, while lighter-weight architectures improve inference efficiency for deployment [Liu2024RoboMamba]. Benchmarks such as LIBERO [Liberoneurips2023] standardize evaluation of compositional generalization and post-deployment robustness.
In this work, we adopt this emerging view of VLAs as frozen backbones: the VLA provides a strong yet imperfect policy, and EFN operates purely as an external correction module without modifying the backbone weights.
@@ 2.2 Post-deployment Learning and Retrieval
A key deployment challenge is improving behavior after a policy has been trained, ideally without repeatedly finetuning large models. Human-in-the-loop and continual-learning frameworks study how robots can collect feedback and adapt online [Liu2024IJRROnTheJob, Liberoneurips2023]. In parallel, retrieval-augmented reinforcement learning associates policies or value functions with an external experience buffer and retrieves similar trajectories at test time to guide decision making [r2a, rarl1, rarl2, rarl3]. In robotics, retrieval has been used to condition policies on past episodes, demonstrations, or cached behaviors for training-free guidance.
On the other hand, residual policy learning improves a strong controller by predicting an additive correction rather than a full action, enabling safer and more sample-efficient finetuning [Johannink2019RPL]. Dense, similarity-shaped rewards and world models [Hafner2025DreamerV3, memory3] further stabilize RL by comparing predicted or observed futures to goal-like targets instead of relying solely on sparse task success.
Our perspective. EFN lies at the intersection of these ideas. It targets the post-deployment regime with a frozen VLA, maintains a live experience bank, retrieves a task-relevant transition in a joint vision–language space, and predicts a residual action that refines the base policy. The residual is trained with dense similarity-shaped rewards that compare the observed next frame to the retrieved successor, combining retrieval-conditioned guidance with residual correction, without ever updating the large backbone.
@@ 2.3 Embodied Reinforcement Learning
Embodied RL trains control policies via interaction with physical or simulated environments, facing sparse rewards and limited data. Prior work addresses these issues through goal relabeling [Andrychowicz2017HER], large-scale real-robot data collection [Bodnar2020Q2Opt], leveraging offline datasets before online improvement [Nair2020AWAC], strong regularization from pixels [Yarats2021DrQv2], and world-model–based imagination [Hafner2025DreamerV3]. Benchmarks for compositional and continual embodied learning further emphasise evaluating robustness and knowledge accumulation [Liberoneurips2023, bench1, bench2, bench3].
Our scope. Rather than training ever larger VLAs, we focus on a lightweight post-deployment mechanism that augments a frozen policy. Concretely, EFN is an Experience Feedback Network trained with Soft Actor–Critic [Haarnoja2018SAC] to map retrieved trajectories and current observations to residual actions that refine the controller, effectively learning an experience-feedback module instead of relearning the VLA.
@@ 3.1 Overview of EFN
EFN wraps a frozen vision–language–action (VLA) policy with an experience-centric controller that improves behavior at deployment without updating the backbone weights. We first design a step-level experience bank and record schema that store both rollout-level instructions and per-step visual keys and actions (Sec. 3.2). Given this bank, EFN learns a residual policy on top of the frozen VLA using a similarity-shaped reward and Soft Actor–Critic, with additional shaping to discourage idling (Sec. 3.3). At deployment, all network parameters are frozen: EFN performs instruction-filtered, efficiency-aware retrieval, applies the trained residual correction at each step, and grows the bank online by appending successful rollouts (Sec. 3.4). Thus, post-deployment adaptation arises purely from memory growth and recall, not from further gradient updates. Figure 1 and Figure 2 summarize EFN in the training and deployment regimes, respectively. We elaborate the rationale behind this design in Appendix C for interested readers.
@@
Figure 3: Visualization of EFN’s language-conditioned retrieval. (a) PCA projection of instruction embeddings $\ell_{\tau}$ grouped by task type: similar instructions form clusters, so EFN can restrict retrieval to a small set of relevant rollouts. (b) PCA projection of experience keys $\mathbf{k}_{i}$ and online queries $\mathbf{q}_{t}$: for a given query, the retrieved top-$k$ neighbors (blue) lie close in this space.
@@ 3.2 Experience Bank Design
Storage schema. We organize experiences by full rollouts $\tau=(s_{1},a_{1},\dots,s_{T},a_{T})$ and insert into the bank every step $(s_{t},a_{t})$ encountered during data collection or deployment, i.e., steps where the robot actually executes a control action. We exclude the initial “blank” waiting steps during simulator loading and robot warm-up from both training and evaluation.
The initial offline bank contains both successful and near-successful trajectories (and occasionally failed rollouts), which we keep because our reward is defined on semantic similarity rather than a binary success flag; see Appendix D for a detailed discussion. During deployment, however, we only append trajectories that reach the goal, using these successful rollouts as high-quality references.
For each rollout $\tau$ we also store a fixed instruction embedding $\ell_{\tau}$ obtained by encoding the task description with the VLA’s language model at the beginning of the episode. At the step level, we record three items: (i) the VLA vision-encoder features $\mathbf{F}_{t}\in\mathbb{R}^{L\times C}$ for frame $s_{t}$ (e.g., token features), (ii) a compact key vector $\mathbf{k}_{t}\in\mathbb{R}^{d_{k}}$ derived from $\mathbf{F}_{t}$ for retrieval, and (iii) the base policy’s raw action $\mathbf{a}_{t}^{(0)}$ executed at that step. The bank therefore stores tuples $\big(\ell_{\tau},\mathbf{F}_{t},\mathbf{k}_{t},\mathbf{a}_{t}^{(0)}\big)$ for all valid $t$ across all trajectories.
@@
At query time, we form a query vector $\mathbf{q}_{t}$ from the current frame via the same fusion, compute cosine similarities $s_{i}=\cos(\mathbf{q}_{t},\mathbf{k}_{i})$ to all keys, and select the top-$k$ indices $\mathcal{N}_{k}(\mathbf{q}_{t})$. We then sample one key from this shortlist with a similarity-biased distribution:
@@
Figure 3 illustrates the structure that EFN exploits in the experience bank: instruction embeddings $\ell_{\tau}$ cluster by task type, making the instruction-filtered candidate set $\mathcal{R}_{n}$ small and focused, and retrieved keys form tight local neighborhoods around each query in the visual space.
@@ 3.3 Learning with Residual Policy Optimization
Problem setup. EFN learns to nudge the base policy by recalling a relevant past experience and adjusting the current action so that the next observation resembles “what happened next” in experience. At step $t$, the inputs are: current visual features $\mathbf{F}_{t}$ and the base policy’s action $\mathbf{a}_{t}^{(0)}$, together with a retrieved experience step $(\hat{\mathbf{F}},\hat{\mathbf{a}},\hat{\mathbf{F}}^{+})$ and its rollout-level instruction embedding $\ell$ (retrieval is defined in the previous subsection). EFN’s actor outputs a residual $\Delta\mathbf{a}_{t}$; the executed control is
@@
Intuitively, $\mathbf{a}_{t}^{(0)}$ preserves the base policy’s competence, and $\Delta\mathbf{a}_{t}$ is an experience-informed correction conditioned on the retrieved transition.
Figure 5 visualizes EFN’s behavior in latent action space. We project both the base actions $\mathbf{a}_{t}^{(0)}$ and EFN-corrected actions $\mathbf{a}_{t}$ to 2D using PCA and draw arrows for the residuals $\Delta\mathbf{a}_{t}$. The arrows form short displacements around the base actions, indicating that EFN acts as a fine-grained correction layer on top of the frozen policy rather than a separate controller that overrides it.
@@
In practice we also regularize the residual magnitude to avoid destabilizing the base behavior:
@@
with $\lambda_{\text{sem}},\lambda_{\text{res}}>0$. This reward directly encodes our training signal without requiring supervised residual labels. Because $r_{t}$ is defined purely from feature similarity rather than a success/failure flag, EFN can be trained on both successful and failed rollouts as long as they contain meaningful state transitions (see Appendix D for discussion).
Figure 4: Reward decomposition along a representative rollout under EFN’s shaped objective. Early in the trajectory the agent makes progress and receives high reward; when it idles near a good view, the lazy penalty suppresses $r_{t}$; once it moves toward the retrieved successor frame again, the progress and motion terms dominate.
SAC objective. We train EFN with Soft Actor–Critic, conditioning both actor and critics on the current and experience context. Let
@@
with temperature $\alpha$ optionally tuned to maintain a target entropy. During training, gradients do not flow through the retrieval targets $\hat{\mathbf{F}},\hat{\mathbf{F}}^{+}$; updates are confined to EFN’s actor, critics, and the context encoder.
Figure 5: Residual corrections in latent action space. We project base actions $\mathbf{a}_{t}^{(0)}$ and EFN-corrected actions $\mathbf{a}_{t}=\mathbf{a}_{t}^{(0)}+\Delta\mathbf{a}_{t}$ to 2D using PCA. Each arrow connects a base action (yellow) to its corrected counterpart (blue), showing that EFN applies small, local residuals $\Delta\mathbf{a}_{t}$ rather than replacing the base policy.
Reward shaping with anti-idling penalty. To retain absolute credit for matching the retrieved next observation while penalizing “similar but no progress” and encouraging shorter rollouts, we define semantic similarity ($\mathrm{sim}\in[0,1]$)
@@
where $w_{\text{abs}},w_{\text{prog}},w_{\text{mot}},w_{\text{lazy}}\!\geq\!0$ and $\lambda_{\text{time}}\!\geq\!0$ are scalars. Equation (13) preserves an absolute similarity bonus ($a_{t}$), adds a progress term ($p_{t}]_{+}$) to credit genuine improvement toward the retrieved next state, encourages non-trivial motion ($m_{t}$) to avoid degenerate idling, and introduces an anti-idling penalty that activates only when the prediction is already similar yet shows negligible improvement and little change between consecutive frames. The per-step cost $\lambda_{\text{time}}$ favors shorter successful trajectories. Conceptually, EFN thus learns a residual policy that steers $\mathbf{F}_{t+1}$ toward $\hat{\mathbf{F}}^{+}$.

### v3

@@ Abstract
Embodied agents face a fundamental limitation: once deployed in real-world environments, they cannot easily acquire new knowledge to improve task performance. In this paper, we propose Dejavu, a general post-deployment learning framework that augments a frozen Vision-Language-Action (VLA) policy with retrieved execution memories through an Experience Feedback Network (EFN). EFN identifies contextually relevant prior action experiences and conditions action prediction on the retrieved guidance. We train EFN with reinforcement learning and semantic similarity rewards, encouraging the predicted actions to align with past behaviors under the current observation. During deployment, EFN continually expands its memory with new trajectories, enabling the agent to exhibit “learning from experience.” Experiments across diverse embodied tasks show that EFN improves adaptability, robustness, and success rates over frozen baselines. Our Project Page is https://dejavu2025.github.io/.
00footnotetext: *Corresponding authors.
@@ 1 Introduction
Embodied intelligence studies agents that learn and act through physical interaction with their environment [37, 26]. Recently, unified Vision–Language–Action (VLA) models have shown impressive generalization across diverse tasks [48, 34, 11, 16], but only after massive offline training on fixed data distributions [6, 40]. Once deployed, their weights, and thus their knowledge, remain fixed unless the model is retrained [27], so most real systems effectively “stop learning” in the wild [27].
@@
This raises a natural question: must intelligent systems always rewrite their internal weights to improve? Humans often tackle new problems by recalling and reusing past experiences rather than acquiring entirely new facts [1, 32]. Episodic memory, which underlies the feeling of “déjà vu,” does not alter core knowledge representations but does enable rapid adaptation by analogy [13]. We therefore ask whether an AI agent can similarly improve by recalling and reusing its own “experiences.” If a neural network could leverage stored memories of situations and solutions to inform its current decisions, it could improve over time at inference simply by accumulating and exploiting new experiences, without gradient-based retraining [4, 33, 22]. This idea of “learning from déjà vu” motivates our approach.
@@
A closely related literature studies retrieval-augmented reinforcement learning and retrieval-augmented embodied agents. Retrieval-augmented [24] RL couples an external experience buffer with a parametric policy or value function, retrieving past trajectories similar to the current state to improve credit assignment or value estimation [13]. In embodied and robotic control, recent methods retrieve prior executions in various forms, such as state–action trajectories or policy snippets, and leverage them during decision making, either by fine-tuning on retrieved data [47, 30] or by conditioning the policy directly on retrieved trajectories as additional context [21, 20, 23]. These works share a retrieval-plus-decision paradigm [13, 47, 30], but they typically target a trainable policy whose weights continue to change during or after deployment [20, 21], operate on static offline corpora rather than live memories that grow during deployment [13, 47], and define retrieval over compact state or task abstractions instead of the rich, open-vocabulary vision–language interface of modern VLAs [47, 23]. As a result, existing retrieval-augmented methods do not offer a simple deployment-time mechanism for continually improving a frozen, unified policy. To our knowledge, EFN is the first post-deployment learning framework for VLA policies that couples a live experience bank with retrieval-conditioned residual correction, enabling improvement through memory growth without backbone finetuning.
@@
To this end, we propose the Experience Feedback Network (EFN), an experience-centric controller for improving pre-trained VLA policies at deployment without modifying their weights. EFN takes the current observation together with an action retrieved from a matched experience and predicts a residual action, which is added to the VLA output to produce the final control. When a strong prior experience exists, EFN learns to exploit it to refine the action; when retrieval is uninformative, it falls back toward the base policy. We train EFN with reinforcement learning using a dense similarity-based reward: the agent receives higher reward when the next observation resembles the next observation in the retrieved experience, providing frequent shaping signals beyond sparse success–failure feedback. Concretely, EFN learns retrieval-conditioned residual control by treating the successor of a retrieved transition as a semantic target for shaping. EFN is optimized with Soft Actor–Critic [14], whose entropy regularization and off-policy sample efficiency facilitate effective reuse of stored experiences.
@@
During deployment, we maintain a live experience bank that is continuously augmented with newly successful rollouts. At each inference step, the agent retrieves a similar trajectory in a joint vision–language embedding space, provides the matched experience action to EFN alongside the current observation, and obtains a residual that refines the base VLA output. This yields a retrieval-augmented yet weight-frozen VLA agent that improves purely by accumulating and reusing its own experiences over time. We integrate EFN with OpenVLA [18], UniVLA [8], and GO-1 [7], and evaluate it in both simulation and on a real-world platform. Extensive experiments show that EFN not only improves success and efficiency over strong VLA baselines, but also outperforms representative retrieval-augmented RL and embodied baselines under the same experience bank and evaluation protocol.
We summarize our main contributions as follows:
@@
  • We propose EFN, to our knowledge the first post-deployment learning framework for frozen VLA policies that improves behavior through online experience accumulation and retrieval-conditioned residual correction, enabling adaptation by memory updates rather than backbone finetuning.
@@
  • We design a live experience bank that stores synchronized vision–language–action trajectories, retrieves task-relevant transitions in a joint embedding space, and trains the residual with similarity-shaped reinforcement signals toward retrieved successors.
@@
  • We integrate EFN with OpenVLA, UniVLA, and GO-1, and benchmark it on LIBERO and a real-world AgiBot-G1 setup, showing consistent gains over retrieval-only, residual-only, retrieval-augmented RL, and test-time training baselines under the same interaction budget.
@@
Figure 1: EFN trains a residual policy with SAC to nudge the base action so the next frame matches the successor in memory.
@@
Figure 2: EFN performs inference by retrieving candidates, applying residual correction, and growing the experience bank online.
@@ 2.1 Vision-Language-Action (VLA) Models
Large-scale VLA policies couple open-vocabulary perception with end-to-end control and achieve strong cross-task generalization in manipulation [35, 45, 44, 9, 6, 48]. Recent open-source generalist policies [40, 18] extend this paradigm to heterogeneous robots and sensors, while lighter-weight architectures improve inference efficiency for deployment [28]. Benchmarks such as LIBERO [25] provide a standardized testbed for compositional generalization and post-deployment robustness.
In this work, we adopt the emerging view of VLAs as frozen backbones: the VLA provides a strong yet imperfect policy, and EFN operates purely as an external correction module without modifying the backbone weights.
@@ 2.2 Post-deployment Learning and Retrieval
A central deployment challenge is improving behavior after a policy has been trained, ideally without repeatedly finetuning large models. Human-in-the-loop and continual-learning frameworks study how robots can collect feedback and adapt online [27, 25]. In parallel, retrieval-augmented reinforcement learning associates policies or value functions with an external experience buffer and retrieves similar trajectories at test time to guide decision making [13, 41, 2, 39]. In robotics, retrieval has been used to condition policies on past episodes, demonstrations, or cached behaviors for training-free guidance.
Residual policy learning improves a strong controller by predicting an additive correction rather than a full action, enabling safer and more sample-efficient finetuning [17, 36]. Dense, similarity-shaped rewards and world models [15, 3] further stabilize RL by comparing predicted or observed futures to goal-like targets instead of relying solely on sparse task success.
Our perspective. EFN lies at the intersection of these ideas. It targets the post-deployment regime with a frozen VLA, maintains a live experience bank, retrieves a task-relevant transition in a joint vision–language space, and predicts a residual action that refines the base policy. The residual is trained with dense similarity-shaped rewards that compare the observed next frame to the retrieved successor, combining retrieval-conditioned guidance with residual correction without ever updating the large backbone.
@@ 2.3 Embodied Reinforcement Learning
Embodied RL trains control policies via interaction with physical or simulated environments, often under sparse rewards and limited data. Prior work addresses these issues through goal relabeling [1], large-scale real-robot data collection [5], leveraging offline datasets before online improvement [31], strong regularization from pixels [43], and world-model–based imagination [15]. Benchmarks for compositional and continual embodied learning further emphasize evaluating robustness and knowledge accumulation [25, 42, 46, 12].
Our scope. Rather than training ever larger VLAs, we focus on a lightweight post-deployment mechanism that augments a frozen policy. Concretely, EFN is an Experience Feedback Network trained with Soft Actor–Critic [14] to map retrieved trajectories and current observations to residual actions that refine the controller, effectively learning an experience-feedback module instead of relearning the VLA.
@@ 3.1 Overview of EFN
EFN wraps a frozen vision–language–action (VLA) policy with an experience-centric controller that improves behavior at deployment without updating the backbone weights. We first design a step-level experience bank and storage schema that record both rollout-level instructions and per-step visual keys and actions (Sec. 3.2). Given this bank, EFN learns a residual policy on top of the frozen VLA using a similarity-shaped reward and Soft Actor–Critic, with additional shaping to discourage idling (Sec. 3.3). At deployment, all network parameters remain frozen: EFN performs instruction-filtered, efficiency-aware retrieval, applies the trained residual correction at each step, and grows the bank online by appending successful rollouts (Sec. 3.4). Post-deployment adaptation therefore arises purely from memory growth and recall, not from further gradient updates. Figure 1 and Figure 2 summarize EFN in the training and deployment pipelines, respectively. We elaborate on the rationale behind this design in Appendix SectionC.
@@
Figure 3: Visualization of EFN’s language-conditioned retrieval. (a) PCA projection of instruction embeddings $\ell_{\tau}$ grouped by task type: similar instructions form clusters, allowing EFN to restrict retrieval to a small set of relevant rollouts. (b) PCA projection of experience keys $\mathbf{k}_{i}$ and online queries $\mathbf{q}_{t}$: for a given query, the retrieved top-$k$ neighbors (blue) lie nearby in this space.
@@ 3.2 Experience Bank Design
Storage schema. We organize experience as full rollouts $\tau=(s_{1},a_{1},\dots,s_{T},a_{T})$ and insert into the bank every step $(s_{t},a_{t})$ encountered during data collection or deployment, i.e., every step at which the robot executes a control action. We exclude the initial “blank” waiting steps during simulator loading and robot warm-up from training and evaluation.
The initial offline bank contains both successful and near-successful trajectories, and occasionally failed rollouts, which we retain because our reward is defined on semantic similarity rather than a binary success flag; see Appendix SectionD.1 for discussion. During deployment, however, we append only trajectories that reach the goal, using these successful rollouts as high-quality references.
For each rollout $\tau$, we also store a fixed instruction embedding $\ell_{\tau}$, obtained by encoding the task description with the VLA’s language model at the beginning of the episode. At the step level, we record three items: (i) the VLA vision-encoder features $\mathbf{F}_{t}\in\mathbb{R}^{L\times C}$ for frame $s_{t}$ (e.g., token features), (ii) a compact key vector $\mathbf{k}_{t}\in\mathbb{R}^{d_{k}}$ derived from $\mathbf{F}_{t}$ for retrieval, and (iii) the base policy’s raw action $\mathbf{a}_{t}^{(0)}$ executed at that step. The bank therefore stores tuples $\big(\ell_{\tau},\mathbf{F}_{t},\mathbf{k}_{t},\mathbf{a}_{t}^{(0)}\big)$ for all valid $t$ across all trajectories.
@@
At query time, we form a query vector $\mathbf{q}_{t}$ from the current frame using the same fusion, compute cosine similarities $s_{i}=\cos(\mathbf{q}_{t},\mathbf{k}_{i})$ to all keys, and select the top-$k$ indices $\mathcal{N}_{k}(\mathbf{q}_{t})$. We then sample one key from this shortlist with a similarity-biased distribution:
@@
Figure 3 illustrates the structure EFN exploits in the experience bank: instruction embeddings $\ell_{\tau}$ cluster by task type, making the instruction-filtered candidate set $\mathcal{R}_{n}$ small and focused, while retrieved keys form tight local neighborhoods around each query in the visual space.
@@ 3.3 Learning with Residual Policy Optimization
Problem setup. EFN learns to nudge the base policy by recalling a relevant past experience and adjusting the current action so that the next observation resembles “what happened next” in that experience. At step $t$, the inputs are the current visual features $\mathbf{F}_{t}$ and the base policy’s action $\mathbf{a}_{t}^{(0)}$, together with a retrieved experience step $(\hat{\mathbf{F}},\hat{\mathbf{a}},\hat{\mathbf{F}}^{+})$ and its rollout-level instruction embedding $\ell$ (retrieval is defined in the previous subsection). EFN’s actor outputs a residual $\Delta\mathbf{a}_{t}$; the executed control is
@@
Intuitively, $\mathbf{a}_{t}^{(0)}$ preserves the base policy’s competence, while $\Delta\mathbf{a}_{t}$ provides an experience-informed correction conditioned on the retrieved transition.
Figure 5 visualizes EFN’s behavior in latent action space. We project both the base actions $\mathbf{a}_{t}^{(0)}$ and EFN-corrected actions $\mathbf{a}_{t}$ to 2D using PCA and draw arrows for the residuals $\Delta\mathbf{a}_{t}$. The arrows form short displacements around the base actions, indicating that EFN acts as a fine-grained correction layer on top of the frozen policy rather than as a separate controller that overrides it.
@@
In practice, we also regularize the residual magnitude to avoid destabilizing the base behavior:
@@
with $\lambda_{\text{sem}},\lambda_{\text{res}}>0$. This reward directly encodes our training signal without requiring supervised residual labels. Because $r_{t}$ is defined purely from feature similarity rather than a success/failure flag, EFN can be trained on both successful and failed rollouts as long as they contain meaningful state transitions.
Figure 4: Reward decomposition along a representative rollout under EFN’s shaped objective. Early in the trajectory, the agent makes progress and receives high reward; when it idles near a good view, the lazy penalty suppresses $r_{t}$; once it moves toward the retrieved successor again, the progress and motion terms dominate.
SAC objective. We train EFN with Soft Actor–Critic, conditioning both actor and critics on the current and retrieved experience context. Let
@@
with temperature $\alpha$ optionally tuned to maintain a target entropy. During training, gradients do not flow through the retrieval targets $\hat{\mathbf{F}}$ and $\hat{\mathbf{F}}^{+}$; updates are confined to EFN’s actor, critics, and the context encoder.
Figure 5: Residual corrections in latent action space. We project base actions $\mathbf{a}_{t}^{(0)}$ and EFN-corrected actions $\mathbf{a}_{t}=\mathbf{a}_{t}^{(0)}+\Delta\mathbf{a}_{t}$ to 2D using PCA. Each arrow connects a base action (yellow) to its corrected counterpart (blue), showing that EFN applies small, local residuals $\Delta\mathbf{a}_{t}$ rather than replacing the base policy outright.
Reward shaping with anti-idling penalty. To retain absolute credit for matching the retrieved next observation while penalizing “similar but no progress” behavior and encouraging shorter rollouts, we define semantic similarity ($\mathrm{sim}\in[0,1]$)
@@
where $w_{\text{abs}},w_{\text{prog}},w_{\text{mot}},w_{\text{lazy}}\!\geq\!0$ and $\lambda_{\text{time}}\!\geq\!0$ are scalars. Equation (13) preserves an absolute similarity bonus ($a_{t}$), adds a progress term ($p_{t}]_{+}$) to credit genuine improvement toward the retrieved next state, encourages non-trivial motion ($m_{t}$) to avoid degenerate idling, and introduces an anti-idling penalty that activates only when the prediction is already similar yet shows negligible improvement and little change between consecutive frames. The per-step cost $\lambda_{\text{time}}$ favors shorter successful trajectories. Conceptually, EFN therefore learns a residual policy that steers $\mathbf{F}_{t+1}$ toward $\hat{\mathbf{F}}^{+}$.